solve for z, then y, then x
-z=3xy+2xz-1

Answer by checkley77(12844)

-z=3xy+2xz-1
3xy+2xz=1-z
x(3y+2z)=1-z
x=(1-z)/(3y+2z) answer.
3xy=1-2xz-z
y=(1-2xz-z)/3x answer.
2xz+z=-3xy+1
z(2x+1)=-3xy+1
z=(-3xy+1)/(2x+1) answer.